SuicidalSheep32:

Each cone of the hourglass has a height of 15 millimeters. The total height of the sand within the top portion of the hourglass is 45 millimeters. The radius of both cylinder and cone is 6 millimeters. Sand drips from the top of the hourglass to the bottom at a rate of 10π cubic millimeters per second. How many seconds will it take until all of the sand has dripped to the bottom of the hourglass?

SuicidalSheep32:

126 108 18 29

SuicidalSheep32:

where do i even start?

Hero:

How did you get from an extremely basic geometry question to an advanced calculus question in a matter of minutes?

SuicidalSheep32:

Honestly my class is confusing

Hero:

I already work for a company where I answer questions on this level and I get paid to write the answers. I'd rather go to that site and answer these kinds of questions than do it for free here. Sorry to disappoint you. I can answer up to basic level calculus, but for something like this, and only because of the amount of time required to answer, I can only devote so much time here.

SuicidalSheep32:

what is the site?

Hero:

Slader.com, but I'm already assigned books to work on. I can't work on the book you're working on unless the question is in one of my already assigned books.

SuicidalSheep32:

huh so what do i need to do

SuicidalSheep32:

never mind the site is blocked on this computer

SuicidalSheep32:

thank you anyway

Hero:

That's not surprising

Hero:

You'd have access to all the answers in your textbook.

SuicidalSheep32:

i just needed to be explained to but thanks bro i might need help with another simple one soon though

Cupcake123456:

Calculate the volume of sand, the cone will be completely filled and the cylinder will have sand up to the 30 mm level. Volume of sand will = the volume of the cone: 1/3 * pi * (6 mm)^2 * 15 = 1/3 * pi * 36 * 15 =180 pi cubic mm the cylinder will have a volume of sand equal to: pi * (6 mm)^2 * 30 mm=pi * 36 sq mm * 30 mm=1080 pi cubic mm. The total sand is the sum: 1080 pi cubic mm +180 pi cubic mm = 1260 pi cubic mm.

Hero:

Yeah I get you, but what you don't understand is that I was supposed to start working six hours ago. I procrastinated thus far and now I really have to start working.

Cupcake123456:

the sand travels at a rate of 10 pi/sec so the sane will take (1260 pi)/10 pi/sec= =126 pi seconds. ... i may have made an error so .... someone should look at this @Hero

SuicidalSheep32:

well i think it makes sense and it looks right but i am not sure either

Hero:

Looks right @Cupcake123456 . I was having trouble imagining what the hourglass would look like. But I realize that the bottom portion would be the cone and the cylinder would be above the cone and all the sand would fill both. I think you have the right idea, but double-check your calculations.

umm:

@Cupcake123456 - You should always post where you get your sources from. >Brainly.

Cupcake123456:

@umm no i didnt use brainly but thank you anyway only sources i used were my brain and a calculator

Hero:

Also, @Cupcake123456 remember this is not an answer-giving site. This is an "explain to the user how to do the problem then let them do the work themselves" site.

Cupcake123456:

you told me dat in my inbox yeen have to come back and post it ...